OpenStudy (11study11):

Help 1 question!!!!!!!!!! multiple choice!!!!!!!!

OpenStudy (11study11):

The table below shows four systems of equations: System 1 System 2 System 3 System 4 4x − 5y = 2 3x − y = 4 4x − 5y = 2 10x − 21y = 10 4x − 5y = 2 24x − 47y = 22 4x − 5y = 2 10x + 3y = 15 Which pair of systems will have the same solution? System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to two times the second equation in system 1 System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to two times the second equation in system 2 System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to three times the second equation in system 1 System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to three times the second equation in system

OpenStudy (11study11):

@zepdrix @Awolflover1 @inkyvoyd Please helpppppp!!!!!!!!!!!!